One number is 4 more than 3/8 of another number. Total of both is 427. Please give formula. Thanks
Answer by nerdybill(7384)

One number is 4 more than 3/8 of another number. Total of both is 427.
.
Let x = one number
and y = another number
.
From: "One number is 4 more than 3/8 of another number."
x = 4+ (3/8)y (equation 1)
.
From: "Total of both is 427"
x + y = 427 (equation 2)
.
Solve equation 2 for y:
x + y = 427
y = 427-x
.
Substitute above into equation 1 to find x:
x = 4+ (3/8)y
x = 4+ (3/8)(427-x)
8x = 32+ (3)(427-x)
8x = 32+ 1281-3x
11x = 32+ 1281
11x = 1313
x = 119.36
.
Substitute above into equation 2 to find y:
x + y = 427
119.36 + y = 427
y = 307.64